OpenStudy (anonymous):

can some one help me i need to know how to graph this f(x)=5x

OpenStudy (anonymous):

plug in values for x, find corresponding y coordinates. plot the points on the graph.

OpenStudy (anonymous):



OpenStudy (anonymous):

that is what the graph will look like but it is not to scale. you have to plot the points properly

OpenStudy (anonymous):

how did you get that ?i dont get it

OpenStudy (anonymous):

ok how did you get the points?

OpenStudy (anonymous):

k lets take 3 points for x. -1,0,1 f(-1)=5(-1)=-5 f(0)=5(0)=0; this is the origin f(1)=5(1)=5 so three points you can use are (-1,5), (0,0), (1,5). connecting those points with a line will show a graph similar to the one above

OpenStudy (anonymous):

you can choose ANY points you want for x. because the domain is all real numbers

OpenStudy (anonymous):

so you can choose any points

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

as long as they are in the domain of the function. for example if f(x) = 1/x; then x would not be allowed to equal zero. so you wouldn't use zero. because you can't divide by zerio.